# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v1

@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques. In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [131]. As shown in Table IV, numerous approaches including PDCO [129], Hua et al. [96], and ETCRA [94] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making have to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [131, 130]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [96], and user mobility patterns [128]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [111], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [135]. Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [96]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [128]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [111]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [135].
@@ IV-B2 Collaborative Resource Management
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [70]. In this dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [156]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [93]. This strategy allows edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [93]. Consequently, this adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhances performance through efficient network-wide resource utilization [118].
@@ V-C Communication Security
Securing communication channels is paramount in edge-cloud collaborative systems to protect data transmission from unauthorized access and tampering. This protection involves secure data transmission methods, including encryption and secure communication protocols, safeguarding data in transit. Authentication and access control mechanisms, such as multi-factor authentication and role-based access control, prevent unauthorized access to resources and data. Various encryption schemes, including symmetric and asymmetric encryption, secure data storage and transmission in edge-cloud environments. Additionally, blockchain technology offers potential for enhancing security by providing decentralized security, data integrity, and transparent audit trails.
@@ VI-B Smart Healthcare
Advancements in AI and edge-cloud computing are transforming smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig.7. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [237]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [238]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [239], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, processed at the edge for immediate feedback and transmitted to the cloud for analysis and storage [240]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [241]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [242]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [243]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [244].
@@ VII-A Evaluation Metrics
Performance evaluation of edge-cloud collaborative systems requires a comprehensive assessment across multiple dimensions. We categorize the relevant metrics into four primary groups: system performance, model efficiency, quality of service, and security and privacy. System performance metrics. Fig. 8 shows comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering of real-world constraints like network variability, compute heterogeneity, and regulatory demands.

### v2

@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques. In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [136]. As shown in Table IV, numerous approaches including PDCO [134], Hua et al. [102], and ETCRA [101] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [136, 135]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [102], and user mobility patterns [133]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [116], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [140]. Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [102]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [133]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [116]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [140].
@@ IV-B2 Collaborative Resource Management
One approach leverages optimal transport (OT) and FL principles. Specifically, an OT-based offline optimization model can initially allocate bandwidth and computation resources, followed by online optimization using federated actor-critic learning (OTFAC) [77]. In this dual-driven approach, edge servers can act as local aggregators within the federated learning process, improving efficiency and reducing delays [160]. Additionally, collaborative resource management benefits from a hybrid approach where edge servers not only collaborate with the cloud but also expand their local resources [100]. This strategy allows edge servers to dynamically purchase cloud resources and expand local capacity for long-term cost optimization [100]. Consequently, this adaptability manages dynamic workloads and varying demands. Edge-cloud and edge-edge cooperation further enhance performance through efficient network-wide resource utilization [123].
@@ V-C Communication Security
Securing communication channels is paramount in edge-cloud collaborative systems to protect data transmission from unauthorized access and tampering. This protection involves secure data transmission methods, including encryption and secure communication protocols, safeguarding data in transit. Authentication and access control mechanisms, such as multi-factor authentication and role-based access control, prevent unauthorized access to resources and data. Various encryption schemes, including symmetric and asymmetric encryption, secure data storage and transmission in edge-cloud environments. Additionally, blockchain technology offers the potential to enhance security by providing decentralized security, data integrity, and transparent audit trails.
@@ VI-B Smart Healthcare
Advancements in AI and edge-cloud computing are transforming smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig.7. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [241]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [242]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [243], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [25]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [244]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [245]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [246]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [247].
@@ VII-A Evaluation Metrics
Performance evaluation of edge-cloud collaborative systems requires a comprehensive assessment across multiple dimensions. We categorize the relevant metrics into four primary groups: system performance, model efficiency, quality of service, and security and privacy. System performance metrics. Fig. 8 shows comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering real-world constraints like network variability, compute heterogeneity, and regulatory demands.